### v3

# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

Abstract
Through this comprehensive survey of Kolmogorov-Arnold Networks (KAN), we have gained a thorough understanding of its theoretical foundation, architectural design, application scenarios, and current research progress. KAN, with its unique architecture and flexible activation functions, excels in handling complex data patterns and nonlinear relationships, demonstrating wide-ranging application potential. While challenges remain, KAN is poised to pave the way for innovative solutions in various fields, potentially revolutionizing how we approach complex computational problems.
1 Introduction
1.1 Research Background
Kolmogorov-Arnold Networks (KAN) are an emerging neural network architecture based on the theorems of Kolmogorov and Arnold. These theorems demonstrate that any continuous multivariable function can be represented as a superposition of a finite number of univariate functions. This discovery provides a solid mathematical foundation for the design of KAN, allowing it to replace the fixed linear weights in traditional neural networks with learnable univariate functions, thereby enhancing the model’s flexibility and interpretability.
Kolmogorov’s theorem, proposed in 1957, states that any continuous multivariable function can be represented as a finite superposition of univariate functions. This research laid the theoretical groundwork for the development of KAN, enabling it to be applied to more complex function representation and learning tasks [1]. The application of KAN is supported by several classical works in numerical analysis and partial differential equation (PDE) solving. For instance, Babuska and Rheinboldt [2] discussed the application of adaptive finite element methods in solving PDEs, providing theoretical background for the application of KAN in numerical analysis. The CFL condition, introduced by Courant, Friedrichs, and Lewy [3], is crucial in numerical analysis and also supports the understanding of KAN in PDE solving. Additionally, the regularization method proposed by Tikhonov [4] plays a key role in dealing with ill-posed problems and PDE solving, significantly influencing the application of KAN in data fitting techniques [2, 3, 4].
In recent years, the application of KAN in machine learning has attracted widespread attention. Its main advantages lie in the flexibility and interpretability of the model. Unlike traditional multilayer perceptrons (MLP), KAN does not have fixed activation functions but uses learnable univariate functions. This characteristic allows KAN to excel in data fitting and complex function learning tasks. For example, Liu et al. [5] demonstrated the superior performance of KAN in data fitting and PDE solving, showing that even smaller KAN models can achieve or surpass the performance of MLPs in these tasks [5].
Another significant advantage of KAN is its performance in handling high-dimensional data. Xu et al. [6] applied KAN in time series analysis, demonstrating KAN’s capability in capturing complex temporal dependencies. Furthermore, the application of KAN in graph-structured data processing and hyperspectral image classification also showcases its broad applicability and excellent performance [7].
1.2 Research Objectives
This paper aims to systematically review the theoretical foundation, implementation methods, and performance of KAN in various applications. The specific objectives include:
  1. 1.
    To elucidate the basic structure and working principles of KAN.
  2. 2.
    To explore the implementation methods and challenges of KAN in different tasks.
  3. 3.
    To analyze the performance of KAN in practical applications and compare it with other models.
  4. 4.
    To summarize the current research progress and propose future research directions.
Through this review, readers will gain a comprehensive understanding of the current state of KAN research and its development potential, providing valuable references for subsequent research and applications.
2 Kolmogorov-Arnold Representation Theorem Overview
2.1 Historical Background and Proponents
The Kolmogorov-Arnold theorem was proposed by the renowned Russian mathematician Andrey Kolmogorov in 1957, addressing the representation problem of continuous multivariable functions. Kolmogorov’s work is considered a milestone in the field of mathematics as it provided a revolutionary method to decompose complex multivariable functions into simple univariate functions. This theoretical breakthrough had a profound impact on approximation theory and laid the groundwork for subsequent research.
Andrey Kolmogorov was one of the most influential mathematicians of the 20th century, making significant contributions to various fields including probability theory, topology, mechanics, and function approximation. In 1957, Kolmogorov published his research on the representation of continuous multivariable functions, proposing that any continuous multivariable function can be represented as a superposition of a finite number of continuous univariate functions [1]. This discovery greatly advanced the fields of function approximation and numerical analysis.
Following Kolmogorov’s theoretical proposal, another prominent Russian mathematician, Vladimir Arnold, further extended Kolmogorov’s work in 1958. Arnold’s research validated the universality of Kolmogorov’s theorem and demonstrated its effectiveness in representing higher-dimensional functions. He proved that even complex high-dimensional functions can be approximated by a finite number of simple univariate functions [8]. Arnold’s work not only reinforced the foundation of Kolmogorov’s theory but also directed further research in this area.
The research contributions of Kolmogorov and Arnold are significant not only in theoretical terms but also in practical applications. By decomposing complex multivariable functions into simple univariate functions, the Kolmogorov-Arnold theorem provides powerful tools and methods for fields such as machine learning, data fitting, and partial differential equation (PDE) solving. The proposal and validation of this theorem mark an important milestone in mathematical analysis and applied mathematics.
2.2 Core Content and Mathematical Expression of the Theorem
Kolmogorov’s theorem states that any continuous multivariable function can be represented as a superposition of a finite number of continuous univariate functions. Specifically, for an n-dimensional continuous function $f(x_{1},x_{2},\ldots,x_{n})$, there exists a set of continuous univariate functions $g_{i}$ and $h_{ij}$ such that:
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{i=1}^{2n+1}g_{i}\left(\sum_{j=1}^{n}h_{ij}(x_{j})\right)$
This theorem indicates that complex multivariable functions can be represented through combinations of simple univariate functions, providing new perspectives for neural network design [1]. This form of representation addresses part of the challenge posed by Hilbert’s thirteenth problem regarding the decomposition of multivariable functions, highlighting the feasibility of function decomposition [8, 9].
In the implementation of Kolmogorov-Arnold Networks (KAN), B-splines (basic spline functions) are used for function approximation. B-splines are constructed based on basic spline functions by selecting appropriate knots and applying the Cox-de Boor recursion formula. This approximation method can effectively represent the local characteristics of complex functions in multiple dimensions while maintaining overall mathematical rigor. KAN utilizes this approach to more accurately model and predict multidimensional data, demonstrating the practical application value of the Kolmogorov-Arnold theorem in modern data science and neural network design.
2.3 Applications in Function Approximation and Multivariable Function Decomposition
Despite the theoretical power of the Kolmogorov-Arnold theorem, there are some limitations in practical applications. For instance, highly discontinuous functions and those with complex topological structures may encounter difficulties when represented using this theorem. Compared to multilayer perceptrons (MLPs), Kolmogorov-Arnold Networks (KAN) theoretically possess greater expressive power, as they can represent arbitrary multivariable functions. However, in practice, KANs have higher computational complexity, especially when a large number of function superpositions are required, making them less efficient than MLPs [10, 11].
Attempts to apply the Kolmogorov-Arnold theorem to neural networks began in the late 20th century. However, due to limited computational capacity and algorithm optimization technologies at the time, these attempts did not achieve significant success. Specifically, the computational resources available then could not effectively handle the large number of function superpositions and complex nonlinear relationships, limiting the practical application of KANs [12, 13].
In recent years, the Kolmogorov-Arnold theorem has regained attention due to the work of Liu et al. They improved the performance of neural networks in handling complex data relationships by introducing learnable univariate functions [5]. Additionally, new research efforts have rapidly emerged in this field, such as Wavelet-KAN (Wav-KAN) and the quantum version of KAN (VQKAN), further expanding the application scope and performance of KANs [10, 9]. Therefore, it is necessary to recompile the related work on KAN networks, providing a comprehensive reference for researchers in this field. This will not only help understand the existing research outcomes but also offer valuable guidance for future research directions.
3 Kolmogorov-Arnold Network Architecture
3.1 Basic Structure and Features of KAN Networks
The architecture of Kolmogorov-Arnold Networks (KAN) is unique in that its core idea is to replace traditional fixed linear weights with learnable univariate functions, achieving greater flexibility and adaptability. The architecture of KAN consists of multiple layers, each containing several nodes and edges. Each node is responsible for receiving input signals from the previous layer and applying nonlinear transformations to these signals via learnable univariate functions on the edges. These univariate functions are typically parameterized by spline functions to ensure smoothness and stability in the data processing process [5].
The advantage of this architecture is its ability to dynamically adapt to different data patterns, enhancing the model’s expressiveness. Specifically, each weight parameter in KAN is represented as a univariate function, adjusted during the training process to adapt to different data patterns and improve the network’s flexibility and accuracy [10, 5].
3.2 What is KAN
Univariate Function Representation
A core feature of KAN is that each weight parameter is represented by a learnable univariate function. These functions are typically implemented using spline functions or other suitable parameterization methods. This design allows KAN to apply dynamically adjusted activation functions on the network edges, rather than using fixed activation functions at the nodes as in traditional neural networks.
Dynamic Activation Functions
In KAN, activation functions are not limited to predefined forms (such as ReLU, sigmoid, etc.) but can be dynamically adjusted during the training process. This flexibility enables KAN to adapt to various complex data patterns and nonlinear relationships.
Architectural Flexibility
The architecture of KAN can be customized and adjusted according to the specific tasks and datasets. This flexibility is reflected in the representation of weight parameters and the dynamic adjustment of activation functions.
3.3 What KAN is Not
Fixed Weights and Activation Functions
Traditional neural networks, such as multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), use fixed linear weights and predefined activation functions. In these networks, the weight parameters are fixed during the training process and do not dynamically adjust.
Static Architectural Design
The architectural design of traditional neural networks is static and does not change during the training process. The forms of weights and activation functions are determined when the network is constructed.
3.4 Methods Unique to KAN
Weight Representation Based on Kolmogorov-Arnold Theorem
KAN uniquely employs the Kolmogorov-Arnold theorem, representing weight parameters through learnable univariate functions. This method allows KAN to capture more complex nonlinear relationships, which is not possible with traditional neural networks.
Dynamically Adjusted Activation Functions
KAN introduces dynamic activation functions, meaning the forms of activation functions can be adjusted based on data requirements during the training process. This flexibility significantly enhances the model’s expressiveness and adaptability.
Extensive Parameter Optimization
Unlike traditional neural networks that only optimize fixed weights and biases, KAN’s optimization process includes optimizing the parameters of activation functions. This multi-level optimization provides greater flexibility, allowing KAN to more accurately capture the complex relationships in data.
3.5 Specific Technical Optimizations
Residual Activation Functions
Using base functions $b(x)$ (similar to residual connections), where $b(x)=\text{silu}(x)=\frac{x}{1+e^{-x}}$, making the activation function $\Phi(x)$ a sum of the base function and spline functions.
Initialization Method
Each activation function is initialized as $spline(x)\approx 0$ (by setting $c_{i}\sim N(0,\sigma^{2})$, typically $\sigma=0.1$), while $w$ is initialized in the same way as the linear layers in MLPs.
Updating the Spline Grid
Dynamically update each grid based on input activation to address the problem of splines being defined on a bounded region while activation values may exceed this fixed region during training.
3.6 Is Deep Learning a Necessary Component of KAN?
The design philosophy of Kolmogorov-Arnold Networks (KAN) is rooted in the Kolmogorov-Arnold theorem, emphasizing the representation of weight parameters through univariate functions. This architectural design is compatible with the fundamental ideas of deep learning, but whether KAN requires all components of deep learning and the relationship between the two needs further exploration.
Necessity
Training Process
KAN relies on deep learning training methods such as gradient descent and backpropagation. These methods are indispensable in KAN’s optimization process, making the training process of deep learning essential for KAN. These optimization algorithms have proven effective in adjusting a large number of parameters, achieving good generalization performance.
Nonlinear Representation Capability
The multi-layer network architecture and nonlinear activation functions in deep learning provide strong nonlinear representation capabilities, consistent with KAN’s goal of capturing complex relationships through dynamically adjusted activation functions. Therefore, the core ideas of deep learning are irreplaceable in enhancing KAN’s performance. Through the multi-layer structure, deep learning can extract abstract features of data layer by layer, which is crucial in handling complex data.
Scalability and Adaptability
The scalability and adaptability of deep learning models are important aspects that KAN can leverage. Deep learning models have demonstrated strong adaptability and scalability in various fields such as computer vision, natural language processing, and reinforcement learning, and these experiences can be directly applied to the research and development of KAN.
Alternatives
Traditional Machine Learning Methods
Although KAN can leverage deep learning training methods, its unique representation of weight parameters through univariate functions and the design of dynamic activation functions differentiate it from traditional deep learning models. Theoretically, other optimization methods (such as evolutionary algorithms or Bayesian optimization) could be used to train KAN, but these methods might not fully realize KAN’s potential. Traditional methods might be less efficient in handling high-dimensional and complex datasets.
Other Architectures
If deep learning methods are not used to train KAN, new algorithms might need to be developed to address its complex optimization problems. However, the effectiveness and efficiency of such methods in practical applications remain to be verified. For example, symbolic regression or genetic algorithms could be used for similar tasks, but their scalability and optimization efficiency might not match that of deep learning techniques.
3.7 Is KAN a Branch of Deep Learning?
Theoretical Foundation
The theoretical foundation of KAN is the Kolmogorov-Arnold theorem, which differs from the multi-layer nonlinear network models of deep learning. However, both aim to capture complex nonlinear relationships through hierarchical structures. Deep learning models achieve this through architectures like multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), while KAN does so through dynamically adjusted activation functions and univariate function weights.
Architectural Features
Although KAN features unique univariate function weight representations, its hierarchical network structure and dynamic activation function design share similarities with neural network models in deep learning. Therefore, KAN can be seen as a special branch of deep learning, aiming to enhance model flexibility and expressiveness through new methods.
Application Domains
KAN’s application domains include data fitting, partial differential equation solving, hyperspectral image classification, and time series analysis, which are traditionally strengths of deep learning. KAN demonstrates the potential to surpass traditional deep learning models in these areas through its unique architecture and optimization methods.
Figure 1: Example of KAN’s 655Application
3.8 Position and Role of Activation Functions in KAN Networks
One of the key features of KAN is the use of learnable activation functions. These activation functions are typically constructed from spline functions, a type of piecewise polynomial function defined by a set of control points that determine their shape. These control points can be adjusted independently, allowing local changes in the function’s shape without affecting its global properties. This characteristic ensures that spline functions maintain the smoothness of the activation functions while flexibly adapting to complex patterns in the data [10].
By using spline functions, KAN is able to apply complex nonlinear transformations on each edge, enhancing the network’s performance in handling nonlinear relationships. Specifically, the diversity and flexibility of spline functions ensure that the model remains efficient and stable when dealing with different types of data [5].
3.9 Function Approximation Capability and Neural Scaling Effect of KAN Networks
Kolmogorov-Arnold Networks (KAN) enhance their ability to handle nonlinear relationships by using spline functions to implement complex nonlinear transformations on each edge. The diversity and flexibility of spline functions ensure that the model remains efficient and stable when dealing with various types of data. KAN’s training process adopts a backpropagation algorithm similar to that of traditional neural networks, but the core difference lies in simultaneously optimizing the learnable activation functions on the edges and the weights between nodes. This requires a fine-tuned adjustment strategy to ensure model stability and convergence.
To improve training efficiency and generalization ability, research suggests using regularization techniques such as Dropout, weight decay, and batch normalization. Additionally, combining advanced optimization algorithms like Adam and stochastic gradient descent can further optimize KAN’s performance when handling large-scale datasets [5].
The scaling perspective of KAN can be summarized as follows:
  1. 1.
    Simple two-layer KANs may provide non-smooth representations, while deeper KANs can achieve smoother function approximation by increasing the number of layers.
  2. 2.
    Even KANs with a limited number of dimensions can effectively approximate target functions, demonstrating the broad approximation capability of KAN.
The neural scaling law describes the phenomenon where model test loss decreases with an increase in the number of model parameters, following the relationship $\ell\propto N^{-\alpha}$, where $\ell$ represents the root mean square error (RMSE) during testing, $N$ represents the number of parameters, and $\alpha$ is the scaling exponent. This indicates that increasing the number of parameters can significantly improve model accuracy.
KAN leverages the precise approximation capability of B-spline curves by fine-graining them over the data domain, achieving more accurate function representation. This strategy shows the unique advantage of KAN compared to multilayer perceptrons (MLPs), which typically do not train on data partitioning. By starting with a KAN with fewer parameters and gradually refining the spline grid to a model with more parameters, KAN can achieve high-precision approximation with lower complexity.
4 Advantages and Performance of KAN Networks
4.1 Performance Comparison with Traditional Multilayer Perceptrons (MLPs)
Kolmogorov-Arnold Networks (KAN) exhibit significant advantages over traditional neural networks such as Multilayer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs) in several aspects. First, KAN’s unique design is based on the Kolmogorov-Arnold theorem, introducing learnable univariate functions instead of fixed linear weights. This design allows KAN to apply dynamically adjustable activation functions on network edges, unlike traditional neural networks that use fixed activation functions at the nodes. Specifically, each weight parameter in KAN is represented by a univariate function, typically parameterized by spline functions. This architecture enhances the network’s flexibility, enabling it to better adapt to different data patterns [5, 7].
In terms of parameter optimization, traditional neural networks focus primarily on fixed weights and biases. KAN expands the scope of optimization by introducing learnable activation functions. This means that during the optimization process, not only are the weights and biases of the network adjusted, but also the parameters of the activation functions. While this increases the complexity of training, it also provides greater flexibility, allowing KAN to more effectively capture complex data relationships [14].
Furthermore, traditional neural networks often require a large number of parameters and computational resources, especially in large CNNs. This is a significant limitation in resource-constrained environments. KAN, by using learnable univariate functions, significantly reduces the number of parameters. For example, Convolutional KANs (Conv-KAN) achieve high accuracy with significantly fewer parameters than traditional CNNs, effectively reducing computational costs [7]. This reduction not only improves computational efficiency but also enhances the model’s scalability, making it more suitable for resource-constrained environments [5].
4.2 Advantages of KAN Networks in Handling High-Dimensional Data
KAN networks exhibit unique advantages when dealing with high-dimensional data. Because KAN uses learnable univariate functions instead of fixed weights, it can more flexibly adapt to the complex structures of high-dimensional data. Studies have shown that KAN achieves higher accuracy and efficiency in applications such as hyperspectral image classification and time series analysis through a more flexible model architecture [5, 6]. This advantage makes KAN particularly effective in large-scale datasets and high-dimensional data processing.
4.3 Enhanced Accuracy and Interpretability of KAN Networks
The design of KAN not only improves model performance but also enhances interpretability. The fixed activation functions of traditional neural networks limit the understanding of the model’s decision-making process. In contrast, KAN uses learnable activation functions, enabling researchers and users to more intuitively understand and interpret the internal mechanisms and decision-making processes of the model. This is particularly important for applications requiring high interpretability, such as medical diagnosis and financial forecasting [7]. Enhanced interpretability helps build trust and transparency in the model’s decisions, especially in critical application areas.
To enhance the interpretability of KAN, the following strategies are introduced in its design:
  1. 1.
    Sparsification: As a preprocessing step, linear weights are replaced by learnable activation functions. The L1 norm of these activation functions, defined as the average amplitude of their $N_{p}$ inputs, is used. However, the L1 norm alone is insufficient for sparsification and additional entropy regularization is required.
  2. 2.
    Visualization: By setting the transparency of the activation function $\Phi_{l,i,j}$ proportional to $\tanh(\beta A_{l,i,j})$ (where $\beta=3$), small-amplitude functions are ignored, allowing focus on important functions, thus enhancing interpretability.
  3. 3.
    Pruning: After training with sparsification penalties, the network is pruned to form a smaller subnetwork by removing all non-essential neurons. This not only optimizes the model structure but also improves interpretability.
  4. 4.
    Symbolification: An interface is provided to set some activation functions to specified symbolic forms, such as $\cos$ or $\log$, by fitting affine parameters to pre-activation and post-activation samples, allowing outputs to approximate specific symbolic function transformations. This step further promotes symbolic expression and understanding of the model’s decision-making process.
Several studies have shown that the architecture of KAN, by using learnable univariate functions, can more accurately capture complex data relationships and significantly improve model accuracy. For instance, Liu et al. demonstrated that KAN outperforms traditional methods in hyperspectral image classification, with notable improvements in accuracy and robustness [5]. Additionally, Dhiman’s research indicated that KAN excels in time series forecasting, accurately capturing complex patterns and nonlinear relationships in the data [7]. Through the comprehensive application of these strategies, KAN not only enhances performance but also achieves significant improvements in interpretability.
5 Training and Optimization of KAN Networks
5.1 Training Process and Challenges of KAN Networks
The training process of Kolmogorov-Arnold Networks (KAN) is similar to that of other neural network models, typically employing the backpropagation algorithm. Initially, network parameters are initialized, followed by forward propagation to compute the output. The loss function is then used to measure the discrepancy between the predicted and actual results, and the backpropagation algorithm calculates the gradient of the loss function with respect to each parameter. Finally, parameters are updated using gradient descent or other optimization algorithms [7]. Since KAN’s activation functions are learnable, the training process requires simultaneous optimization of both the activation functions on the edges and the weights between nodes. This process necessitates fine-tuned adjustments to ensure stability and convergence [14].
5.2 Techniques and Strategies Explored by Researchers During Training
To enhance the training efficiency and generalization ability of KAN, researchers have explored various techniques and strategies during the training process. For example, using regularization techniques such as Dropout, weight decay, and batch normalization can effectively improve KAN’s training efficiency and generalization ability [5]. Dropout technique randomly discards a portion of neurons to prevent overfitting; weight decay controls model complexity by adding the L2 norm of weights to the loss function; and batch normalization standardizes inputs, improving the stability and speed of model training [7].
Moreover, recent studies have shown that combining advanced optimization algorithms, such as the Adam optimizer and stochastic gradient descent, KAN performs excellently when handling large-scale datasets [14]. The Adam optimizer combines momentum and adaptive learning rates to converge faster to the optimal solution, while stochastic gradient descent updates each mini-batch of data, significantly enhancing training efficiency [5].
5.3 Methods to Ensure Stability and Efficiency of the Training Process
  1. 1.
    Using Regularization Techniques: Techniques such as Dropout and weight decay can prevent overfitting and improve the generalization ability of the model [7].
  2. 2.
    Batch Normalization: Helps accelerate the training process and makes the model more stable during training [5].
  3. 3.
    Optimization Algorithms: Using advanced optimization algorithms, such as the Adam optimizer and stochastic gradient descent, can significantly enhance training efficiency and model performance [14].
6 Applications of KAN Networks
6.1 Symbolic Regression
Symbolic Regression is a method that seeks mathematical formulas or expressions that best represent the data. Unlike traditional regression analysis, symbolic regression not only fits data points but also attempts to discover the underlying mathematical relationships that generate the data. This method is particularly important in scientific research because it can reveal the physical laws and mathematical principles behind the data.
Despite its theoretical potential, symbolic regression faces many challenges in practical applications. Firstly, symbolic regression requires searching for the best mathematical expression in a vast space of formulas, which is computationally intensive. Traditional symbolic regression algorithms, such as those based on genetic programming, can generate highly interpretable models but are often limited by computational resources, making it difficult to handle large-scale datasets and high-dimensional data [15].
Secondly, symbolic regression performs poorly when dealing with noisy data. Real-world data often contains various noise and outliers, making it more difficult to accurately extract the underlying mathematical relationships. Traditional methods tend to overfit or fail to find robust formulas under such circumstances [16].
6.1.1 Advantages of KAN in Symbolic Regression
Kolmogorov-Arnold Networks (KAN) exhibit significant advantages in the application of symbolic regression, particularly in addressing the aforementioned challenges. KAN’s unique architecture and flexible activation functions enable it to effectively perform symbolic regression tasks, as demonstrated below:
High Precision and Interpretability
KAN uses spline functions as activation functions, allowing for more precise fitting of complex nonlinear relationships. Compared to traditional neural networks, KAN can dynamically adjust its activation functions to better capture complex patterns in the data. For example, Liu et al. (2024) demonstrated KAN’s superior performance in symbolic regression, effectively modeling complex data and achieving high-precision formula discovery [5].
Innovations and Improvements
Path Signature Integration: Path signatures are tools that capture the geometric features of time series data. By integrating path signatures, KAN can better understand and predict complex time series patterns. Path signatures effectively capture the geometric shape and structural information of data, providing a more comprehensive perspective for modeling time series data [17].
Dynamic Adaptation: KAN can dynamically adapt to different data patterns by learning adjustable parameters of path signatures, improving its performance in multi-task time series forecasting. The introduction of path signatures allows KAN to dynamically adjust according to different time series characteristics, enhancing the model’s adaptability and robustness [18].
The emergence of KAN has had a profound impact on the field of symbolic regression. Firstly, the high precision and interpretability of KAN allow researchers to better understand the mathematical relationships behind the data, thereby advancing scientific discovery. For example, Vaca-Rubio et al. (2024) demonstrated KAN’s superior performance in symbolic regression, showing through experiments that KAN is more accurate and efficient than traditional methods when handling complex datasets. Secondly, the flexibility and robustness of KAN make it particularly effective in dealing with noisy data. By dynamically adjusting activation functions, KAN can accurately identify the symbolic formulas that generate the data even in high-noise environments, providing higher reliability and stability for practical applications [5].
6.2 Time Series Prediction
Kolmogorov-Arnold Networks (KAN) have shown significant potential in time series prediction. Time series prediction involves using past data to forecast future values and has wide applications in fields such as financial market analysis, weather forecasting, and medical monitoring. KAN, with its unique architecture and flexible activation functions, can capture complex patterns and both short-term and long-term dependencies in time series data.
6.2.1 T-KAN and MT-KAN
T-KAN (Time-Kolmogorov-Arnold Network)
The T-KAN model is specifically designed for handling time series data. It leverages the flexibility of KAN by using learnable univariate activation functions to capture nonlinear relationships and complex patterns in time series. Compared to the basic KAN model, T-KAN achieves higher performance through the following improvements:
  1. 1.
    Dynamic Activation Functions: Traditional neural networks typically use fixed activation functions, such as ReLU or Sigmoid, while T-KAN uses learnable univariate activation functions. This allows the model to dynamically adjust its activation patterns, better adapting to variations in time series data [5].
  2. 2.
    Capturing Nonlinear Relationships: T-KAN is more effective at capturing nonlinear relationships in time series data, excelling in tasks such as financial market prediction and energy consumption forecasting. For instance, Xu et al. (2024) demonstrated T-KAN’s superior performance in stock market prediction by effectively modeling historical price and trading volume data, achieving accurate short-term price forecasts.
The innovation of T-KAN lies in its ability to better adapt to nonlinear changes in time series data through learnable activation functions. This flexibility makes T-KAN particularly effective at capturing complex temporal dependencies. Additionally, compared to the basic KAN model, T-KAN optimizes the learning mechanism of activation functions, enabling more precise reflection of dynamic changes in time series data.
MT-KAN (Multi-Task Kolmogorov-Arnold Network)
The MT-KAN model extends the capabilities of T-KAN to multi-task learning, enabling it to handle multiple related tasks simultaneously. Compared to the basic KAN model and T-KAN, MT-KAN achieves improvements through the following methods:
  1. 1.
    Shared Network Structure: MT-KAN enhances overall prediction performance by sharing network layers and parameters, leveraging inter-task relationships. For example, in multi-task time series forecasting tasks such as power load forecasting and air quality monitoring, MT-KAN achieves higher prediction accuracy by modeling the interactions in multivariate time series.
  2. 2.
    Parameter Sharing and Task Interaction: By sharing certain model parameters, MT-KAN can transfer useful feature representations between tasks, improving the prediction performance of each task. This approach not only enhances model accuracy but also reduces the amount of training data required.
6.2.2 SigKAN
SigKAN (Signature-Weighted Kolmogorov-Arnold Network)
SigKAN enhances KAN’s performance in time series prediction by incorporating path signatures. Path signatures capture the essential geometric features of time series paths. By integrating this information with KAN’s output, SigKAN can provide a more comprehensive representation of temporal data. Specifically:
  1. 1.
    Path Signature Integration: Path signatures are tools that capture the geometric features of time series data. By integrating path signatures, SigKAN can better understand and predict complex time series patterns.
  2. 2.
    Dynamic Adaptation: SigKAN dynamically adapts to different data patterns by learning adjustable parameters of path signatures, improving its performance in multi-task time series forecasting. Inzirillo and Genet (2024) suggested that SigKAN outperforms traditional methods in multi-task time series forecasting [17].
6.2.3 Experiments and Results
In several experiments, T-KAN, MT-KAN, and SigKAN demonstrated their excellent performance in time series prediction. For example, experiments on financial time series data showed that T-KAN and MT-KAN excel in predicting stock implied volatility, significantly outperforming traditional MLP and RNN models. These results indicate that KAN models can achieve high-precision forecasts even in complex and highly volatile financial data environments. SigKAN, by combining the advantages of path signatures and KAN, further enhances performance in multi-task time series forecasting, showing superior performance in tasks such as market trading volume prediction and absolute return prediction.
Research by Vaca-Rubio et al. (2024) also demonstrated the superior performance of KAN in satellite traffic prediction. Experiments showed that KAN models outperform traditional MLP models when dealing with complex satellite traffic data. KAN not only surpasses MLP in prediction accuracy but also demonstrates excellent parameter efficiency, achieving higher prediction performance with fewer trainable parameters [18].
KAN has shown significant advantages in time series prediction, effectively capturing complex temporal dependencies and nonlinear patterns through its unique architecture and flexible activation functions. These characteristics give KAN broad application potential and excellent performance in fields such as financial forecasting and energy management.
6.3 Graph-Structured Data Processing
Kolmogorov-Arnold Networks (KAN) have demonstrated significant advantages in processing graph-structured data, particularly in the fields of graph learning and Graph Neural Networks (GNNs). Graph-structured data is prevalent in various domains such as social networks, molecular structures, and biological networks. Processing these data requires powerful models to capture complex relationships and structural information between nodes. Graph-structured data processing faces several challenges, including:
  1. 1.
    Complex Node Relationships: The relationships and dependencies between nodes and edges in graph-structured data are complex, making it difficult for traditional neural networks to handle these structures directly.
  2. 2.
    High-Dimensional and Large-Scale Graphs: Graph-structured data often have high-dimensional features and large-scale structures, imposing high demands on computational resources and model scalability.
  3. 3.
    Heterogeneity and Diversity: The structure and features of different graphs can vary significantly, requiring models to have strong generalization capabilities to adapt to various graph data.
6.3.1 Graph Kolmogorov-Arnold Networks (GKAN)
GKAN extends the principles of KAN to graph-structured data. By introducing learnable univariate functions on the edges of the graph, GKAN innovatively changes the way information is processed in graph structures, replacing the fixed convolution architecture used in traditional GCNs.
GKAN: Graph Kolmogorov-Arnold Networks
GKAN’s architecture demonstrates innovation and flexibility when processing graph-structured data. Its main features include:
  1. 1.
    Node Feature Aggregation: GKAN employs two main architectures differing in the sequence of aggregation and activation function application. The first architecture aggregates node features (e.g., summation or averaging) before processing through the KAN layer. This method emphasizes applying learnable activation functions after aggregation, allowing the model to better capture complex relationships within local graph structures [19].
  2. 2.
    Embedding Layer Processing: The second architecture applies KAN layers between node embeddings at each layer before aggregation. By introducing learnable univariate activation functions between layers, GKAN can flexibly adapt to feature changes across different graph structures. This approach allows the model to dynamically adjust its representational capacity between different layers, enhancing its effectiveness in processing complex graph data [19].
  3. 3.
    Dynamic Adaptability: GKAN can dynamically adjust according to the structure and features of different graphs, enhancing its adaptability and robustness. This characteristic makes GKAN particularly effective in handling heterogeneous and diverse graph data. Compared to the basic KAN model, GKAN exhibits greater scalability and adaptability when processing large-scale and high-dimensional graph-structured data [5, 18].
Experiments and Results
In multiple experiments, GKAN demonstrated excellent performance in processing graph-structured data. For example, experiments on the Cora dataset showed that GKAN significantly outperforms traditional GCN models in node classification tasks. Specifically, with 100 features, the accuracy of traditional GCN is 53.5
Additionally, GKAN’s performance on other graph datasets also indicates its superiority in handling large-scale and high-dimensional graph-structured data, especially in tasks such as social network analysis and molecular graph analysis, demonstrating its outstanding performance and broad application potential.
6.4 Hyperspectral Image Classification
Kolmogorov-Arnold Networks (KAN) have demonstrated significant advantages in hyperspectral image (HSI) classification, particularly in handling high-dimensional data and complex spatial-spectral correlations. Hyperspectral images contain rich spectral information, which can be used to distinguish different materials, making them widely applicable in fields such as agriculture, environmental monitoring, and military applications. Hyperspectral image classification faces several challenges, including:
  1. 1.
    High-Dimensional Data: Hyperspectral images typically contain hundreds of spectral bands, requiring effective dimensionality reduction and feature extraction methods [20].
  2. 2.
    Complex Spatial-Spectral Correlations: Each pixel in a hyperspectral image contains both spectral and spatial information, increasing the complexity of classification [21].
  3. 3.
    Data Sparsity: In many applications, annotated hyperspectral image data are very limited, making effective classification with small sample sizes a significant challenge [20].
6.4.1 Advantages of KAN in Hyperspectral Image Classification
Kolmogorov-Arnold Networks (KAN) can effectively address these challenges in hyperspectral image classification through their unique architecture and flexible activation functions. Specifically:
Wavelet-based KAN (Wav-KAN)
Wav-KAN introduces wavelet functions as learnable activation functions into the KAN architecture to achieve nonlinear mapping of input spectral features. This method can effectively capture multi-scale spatial-spectral patterns through dilation and translation, significantly enhancing classification performance. For example, Seydi et al. (2024) demonstrated Wav-KAN’s superior performance on benchmark datasets such as Salinas, Pavia, and Indian Pines, with significant improvements in classification accuracy compared to traditional MLP and Spline-KAN models [20].
SpectralKAN
SpectralKAN uses B-spline functions as activation functions, combining these functions to represent multivariate continuous functions for feature extraction and classification of hyperspectral images. SpectralKAN introduces a spatial-spectral KAN encoder, where the spatial KAN encoder extracts spatial features and compresses the spatial dimensions, while the spectral KAN encoder extracts spectral features and classifies them into changing and non-changing categories. This method not only improves classification accuracy but also reduces model parameters, FLOPs, GPU memory, and training and testing time, enhancing computational efficiency [21].
Detailed Analysis
  1. 1.
    High-Dimensional Data Processing: Wav-KAN and SpectralKAN effectively handle the high-dimensional data problem in hyperspectral images by using learnable activation functions. Wav-KAN uses wavelet functions to perform multi-scale decomposition of spectral data, capturing features at different scales, enabling the classification model to comprehensively understand the spatial and spectral information in the data [20].
  2. 2.
    Spatial-Spectral Correlations: SpectralKAN captures complex spatial-spectral correlations in hyperspectral images by introducing a spatial-spectral KAN encoder that separately processes spatial and spectral features. The spatial KAN encoder first extracts spatial features and performs dimensionality reduction, and then the spectral KAN encoder further extracts spectral features and classifies them. This two-stage processing method significantly improves the model’s ability to capture complex spatial-spectral patterns [21].
  3. 3.
    Data Sparsity Issue: By reducing model parameters and improving computational efficiency, SpectralKAN demonstrates strong adaptability in handling small sample hyperspectral image data. Experimental results show that SpectralKAN outperforms traditional methods on multiple datasets, maintaining high classification accuracy even with limited computational resources [21].
Experiments and Results
In multiple experiments, Wav-KAN and SpectralKAN demonstrated excellent performance in hyperspectral image classification. For example, Wav-KAN’s experimental results on the Salinas dataset showed that this method can better capture complex spatial-spectral correlations, significantly improving classification accuracy [20]. SpectralKAN’s experimental results on multiple HSI-CD datasets also showed that this method significantly reduces computational resource requirements while maintaining high classification accuracy [21].
6.5 Quantum Architecture Search
Kolmogorov-Arnold Networks (KAN) have demonstrated significant potential in Quantum Architecture Search (QAS), which aims to optimize circuit design in quantum computing to enhance the efficiency and performance of quantum algorithms. By leveraging KAN’s flexible architecture and interpretability, researchers can capture and optimize key patterns and relationships in quantum computing.
6.5.1 VQ-KAN (Quantum Variational KAN)
The VQ-KAN model is specifically designed for quantum computing optimization problems, utilizing KAN’s flexibility by employing learnable univariate activation functions to capture and optimize the complex relationships between quantum states and operations. Compared to traditional quantum algorithm optimization methods, VQ-KAN improves performance through the following innovations:
  1. 1.
    Dynamic Activation Functions: Traditional quantum optimization algorithms rely on fixed parameters and strategies, whereas VQ-KAN uses learnable univariate activation functions that can dynamically adjust their activation patterns to better adapt to different quantum operations and states [22].
  2. 2.
    Capturing Nonlinear Relationships: VQ-KAN is more effective at capturing nonlinear relationships in quantum operations, excelling in complex quantum circuit optimization tasks. Research shows that VQ-KAN has significant advantages in optimizing the performance and resource utilization of quantum algorithms [22].
Quantum State Representation
By introducing path signatures, VQ-KAN further enhances its performance in quantum state representation. Path signatures capture the geometric features of quantum states, and by combining this information with KAN’s output, VQ-KAN can represent quantum states more comprehensively. This method not only improves representation accuracy but also enhances the model’s robustness and adaptability when handling different quantum states and operations [22].
Experiments and Results
In multiple experiments, VQ-KAN demonstrated excellent performance in quantum architecture search. For example, experiments in quantum circuit optimization showed that VQ-KAN could better capture and optimize complex quantum operations, improving the performance and resource utilization of quantum algorithms [22].
6.6 Medical Image Segmentation and Generation
Kolmogorov-Arnold Networks (KAN) have shown significant advantages in medical image segmentation and generation, particularly in improving segmentation accuracy and generating high-quality images. KAN, through its unique architecture and flexible activation functions, can capture complex spatial-spectral correlations, optimizing medical image processing tasks.
6.6.1 Application of U-KAN in Medical Image Segmentation
U-KAN integrates KAN layers into the U-Net architecture, enhancing the accuracy and interpretability of segmentation models. Specifically:
  1. 1.
    U-KAN (U-Net with Kolmogorov-Arnold Network): U-KAN integrates KAN layers into the U-Net architecture, using learnable univariate activation functions to replace traditional fixed activation functions. This method captures complex nonlinear patterns in input features, improving segmentation model performance. For example, Li et al. (2024) demonstrated U-KAN’s superior performance in medical image segmentation, achieving high accuracy in various medical image segmentation benchmarks [23].
  2. 2.
    U-KAN in Generation Tasks: U-KAN is also applied in noise prediction for diffusion models, improving high-quality image generation by enhancing the U-Net noise predictor. This method shows U-KAN’s potential in generation tasks, capable of producing high-quality task-oriented model architectures [23].
Detailed Analysis
  1. 1.
    Capturing Complex Nonlinear Patterns: U-KAN, through its flexible activation functions, can effectively capture and represent complex nonlinear patterns in medical images, which is crucial for improving image segmentation accuracy and reliability [23].
  2. 2.
    Model Interpretability: Since KAN’s activation functions are learnable, U-KAN models are more interpretable. This allows medical professionals to better understand and trust segmentation results, especially in critical medical diagnoses [23].
Experiments and Results
In multiple experiments, U-KAN demonstrated excellent performance in medical image segmentation and generation. For example, experimental results on various medical image segmentation benchmarks showed that U-KAN significantly improved segmentation accuracy and remained efficient even with limited computational resources [23].
6.7 Word-Level Explainable Language Model
Kolmogorov-Arnold Networks (KAN) have shown potential in natural language processing, particularly in word-level explainable language models. These models aim to provide more transparent and interpretable language models by explaining the meaning and relationships of each word in a sentence.
6.7.1 UKAN (Understandable KAN)
UKAN leverages KAN’s interpretability to construct an explainable language model, where specific neurons encode individual words, and neuron activations can be fully interpreted as the basis of words. To this end, UKAN proposes a continuous word2vec model, where the meaning of a word is represented by a continuous spectrum of its distance from a word base.
  1. 1.
    Continuous Word2vec Model: By representing word meaning as a continuous spectrum of its distance from words in the word base, UKAN provides finer-grained and interpretable word representations in natural language processing [22].
  2. 2.
    Word-Level Explanation: In the UKAN model, the activation of each word can be fully explained by its position in the word base. This method not only enhances model interpretability but also makes the language model’s prediction and decision-making process more transparent and intuitive [22].
Symbolic Regression and Logical Programming
UKAN combines symbolic regression and logical programming techniques to achieve higher interpretability and accuracy in processing complex semantic relationships and reasoning tasks. For example, by utilizing symbolic regression techniques, UKAN can dynamically adjust activation functions to better capture nonlinear relationships between words, enhancing the model’s performance in natural language understanding tasks [22].
Experiments and Results
In multiple experiments, UKAN demonstrated excellent performance in natural language processing tasks. For example, experimental results on language model tasks showed that UKAN’s word-level explainable language model provided higher interpretability and transparency, significantly improving prediction accuracy and user trust [22].
7 Current Research Progress
7.1 Review of Recent Research
Kolmogorov-Arnold Networks (KAN) have achieved significant progress in various fields. Below are some recent important studies:
Fundamental Research and Extensions
Liu et al. (2024) proposed the foundational KAN framework, using learnable univariate activation functions to replace fixed activation functions in traditional neural networks. The study shows that KAN outperforms traditional Multilayer Perceptrons (MLP) in terms of accuracy and interpretability, particularly excelling in data fitting and Partial Differential Equation (PDE) solving tasks [5, 24].
Symbolic Regression
Liu et al. (2024) explored the application of KAN in symbolic regression tasks, demonstrating its advantages in discovering the mathematical relationships and physical laws underlying data. Experimental results indicate that KAN can accurately identify the symbolic formulas generating the data even in high-noise environments [5, 6].
Time Series Prediction
Xu et al. (2024) investigated the application of KAN in time series prediction, proposing two variants: T-KAN and MT-KAN. T-KAN is designed to detect concept drift in time series and explain predictions through symbolic regression of the nonlinear relationships between successive time steps. MT-KAN improves prediction performance by effectively leveraging complex relationships in multivariate time series. Experimental results show that these methods significantly outperform traditional approaches in terms of prediction accuracy and model interpretability [6, 18].
Graph-Structured Data Processing
Liu et al. (2024) introduced GKAN (Graph Kolmogorov-Arnold Networks), extending the principles of KAN to graph-structured data. By introducing learnable univariate functions on graph edges, GKAN innovatively alters the way information is processed in graph structures, replacing traditional convolution architectures. Experimental results indicate that GKAN significantly outperforms traditional GCN models in graph classification and node classification tasks [5, 25].
Wavelet-KAN (Wav-KAN)
Bozorgasl et al. (2024) proposed Wav-KAN, which enhances the interpretability and performance of KAN by incorporating wavelet functions into the KAN framework. Wav-KAN leverages wavelet transforms to capture high-frequency and low-frequency components of the input data, improving classification accuracy, training speed, and robustness, especially in high-dimensional data and complex patterns [25, 20].
7.2 Improvements and Optimizations
Recent years have seen continuous improvements and optimizations in KAN research. Below are some important advancements:
Rational KAN (rKAN)
Afzalaghaei et al. (2024) proposed rKAN, which models using rational function bases. This method employs Padé approximations and rational Jacobi functions, significantly improving KAN’s performance in regression and classification tasks. rKAN enhances model accuracy and computational efficiency through improved activation functions and more efficient parameter update mechanisms [26, 25].
Convolutional KAN (Conv-KAN)
Conv-KAN applies the principles of KAN to Convolutional Neural Networks by introducing learnable univariate functions on convolutional kernels. This approach improves the structure of traditional convolutional layers and excels in image recognition and processing tasks, particularly in handling complex spatial patterns and high-dimensional image data [25, 6].
KAN-Based Adversarial Training (KAT)
Huang et al. (2024) proposed KAT, utilizing the mathematical properties of KAN to enhance the robustness of deep neural networks through adversarial training. Experimental results indicate that KAT significantly improves model robustness against adversarial samples [27, 18].
KAN in Quantum Architecture Search (KANQAS)
Recent studies have explored the application of KAN in Quantum Architecture Search, demonstrating its potential in quantum computing and optimization problems. KANQAS optimizes quantum circuit design, significantly enhancing the efficiency and performance of quantum algorithms [5, 18].
These recent research advancements and optimization methods showcase the broad potential and exceptional performance of KAN in various application fields. Through continuous optimization and expansion, KAN is poised to play an increasingly important role in deep learning and data analysis.
8 Future Research Directions
8.1 Potential Improvements
Structural Optimization
Further optimization of KAN structures is an important direction for future research. Existing studies have shown that learnable univariate activation functions significantly enhance model flexibility and performance. However, there is still much room for optimization. For instance, Afzalaghaei and Kiamari (2024) proposed improving activation function representation capabilities by introducing rational function bases (rKAN), thereby enhancing model accuracy and computational efficiency in regression and classification tasks [26]. Future research can further explore the introduction of other function bases and how to adaptively select the best activation functions for different tasks.
Training Method Improvements
Improving training methods will also be a key research direction. Current research shows that adversarial training (KAT) can significantly enhance model robustness [27]. Future research can explore more training strategies, such as hybrid adversarial training, transfer learning, and self-supervised learning, to further improve model performance in handling high-noise data and small sample data.
Parallel Computing and Hardware Acceleration
Given the high computational complexity of KAN, future research could focus on parallel computing and hardware acceleration technologies. For example, utilizing Graphics Processing Units (GPUs) and Tensor Processing Units (TPUs) for model acceleration, or designing dedicated hardware architectures to enhance KAN’s computational efficiency.
8.2 New Application Scenarios
Medical Diagnosis
The successful application of KAN in hyperspectral image classification indicates its great potential in medical diagnosis. For instance, KAN can be applied to medical image analysis, improving disease detection and diagnosis accuracy by capturing complex spatial-spectral patterns [20].
Financial Prediction
KAN’s ability to capture nonlinear relationships provides broad application prospects in financial market prediction. Future research can explore KAN’s application in stock price prediction, risk management, and quantitative trading, modeling complex market dynamics to enhance prediction accuracy and decision support capabilities [6].
Environmental Monitoring
Hyperspectral imaging has important applications in environmental monitoring, such as land use classification, water quality monitoring, and weather forecasting. KAN’s advantages in handling high-dimensional and complex data make it an ideal tool for environmental monitoring [18].
8.3 Interdisciplinary Combinations and Research Opportunities
Integration with Quantum Computing
Quantum computing is emerging as a new research hotspot. The combination of KAN and quantum computing (KANQAS) promises to significantly enhance the efficiency and performance of quantum algorithms. Future research can explore how to utilize the powerful computational capabilities of quantum computing to further optimize KAN’s training and inference processes [5].
Integration with Bioinformatics
Bioinformatics data are complex and high-dimensional. KAN’s flexibility and strong modeling capabilities hold great potential in genomics, protein structure prediction, and biological network analysis. Future research can explore KAN’s application in bioinformatics, advancing research in the life sciences [5].
Integration with Physics
KAN’s application in symbolic regression demonstrates its potential in the field of physics. Future research can further explore KAN’s application in discovering physical laws, simulating physical processes, and optimizing experimental designs, driving innovation in physics research [5, 6].
9 Conclusion
Through a comprehensive review of Kolmogorov-Arnold Networks (KAN), we have gained a thorough understanding of its theoretical foundations, architectural design, application scenarios, and current research progress. KAN, with its unique architecture and flexible activation functions, excels in handling complex data patterns and nonlinear relationships, showcasing broad application potential.
Firstly, KAN’s theoretical foundation, derived from the theorems of Kolmogorov and Arnold, endows KAN with exceptional performance in data fitting and solving partial differential equations (PDEs). Secondly, KAN’s architecture, by introducing learnable univariate functions to replace fixed activation functions in traditional neural networks, enhances model flexibility and adaptability while also improving interpretability. This makes KAN particularly valuable in fields requiring high interpretability, such as medical diagnosis and financial prediction.
In specific applications, KAN has demonstrated significant advantages in symbolic regression, time series prediction, graph-structured data processing, and hyperspectral image classification. Current research advancements show that researchers are continuously optimizing KAN’s architecture and training methods, improving its performance in various tasks. Future research directions include further optimizing KAN’s structure and training methods, exploring its potential in new application scenarios such as medical diagnosis, financial prediction, and environmental monitoring, and investigating its interdisciplinary combinations with fields like quantum computing, bioinformatics, and physics.
These research directions will not only further KAN’s development but also bring new breakthroughs to the fields of deep learning and data analysis. In summary, as an emerging neural network architecture, KAN exhibits tremendous research and application potential through its theoretical foundation, flexible architecture, and broad application prospects. With continuous optimization and expansion, KAN is expected to play an increasingly important role in future scientific research and practical applications.